OpenStudy (anonymous):

4x^2+8x+3=0 use the quadratic formula to solve the equation

hero (hero):

Myininaya, you better do this one. If I do it, I'm not going to use the quad form.

hero (hero):

4x^2+6x+2x + 3 = 0 2x(2x+3)+1(2x+3) = 0 (2x+3)(2x+1) = 0 x = -3/2, -1/2

hero (hero):

Didn't really need the quadratic formula for this one.